if you need 50lbs of an alloy that is 50% nickel how many pounds of an alloy with 80% must be mixed with an alloy that is 40% nickel
Answer by oberobic(2304)

With problems like this, you have to keep track of how much "pure" stuff you need.
It says you need 50 lb of 50% nickel.
50*.5 = 25 lbs of pure nickel
.
The total of 50 lb is divided between the amount of 40% nickel and 80% nickel
x = the amount of 40% nickel
(50-x) = the amount of 80% nickel
.
.4x + .8(50-x)=25
.4x + 40 - .8x = 25
-.4x = -15
x = 37.5 lb of 40% alloy
.
50-x = 12.5 lb of 80% alloy
.
Check your answer.
.4*37.5 = 15
.8*12.5 = 10
15+10 = 25
.
Done.